Weekly meeting on Xboxes

Europe/Zurich
18/3-008 - CLIC Meeting room (CERN)

18/3-008 - CLIC Meeting room

CERN

20
Show room on map
Description

https://cern.zoom.us/j/6910405761

 

Videoconference
Weekly meeting on Xboxes
Zoom Meeting ID
64847033219
Host
Nuria Catalan Lasheras
Alternative host
Marçà Boronat
Useful links
Join via phone
Zoom URL

Xboxes weekly meeting. 21.02.2024

Attending: Paz Alonso Arias (PAA), Nuria Catalan Las Heras (NClH), Matteo Volpi (MV), Walter Wuensch (WW), Alexei Grudiev (AG), Igor Syratchev (IS), Anisullah Baig (AB), Ping Wang (PW), Riccardo Boldrini (RB), Zaib Un Nisa (ZN), Lee Millar (LM)

Melbox. MV presents last info on Melbox. Everything is ready to start pulsing:

·         Vacuum in line (short line, half a day to reach 10-9 mbar) ok

·         Interlocks, chillers ok

·         Tuned pulse compressor ok

Ready for condition next week. Right now, same configuration as Xbox3 without structures.

NClH asks if loads are conditioned. Answer: should be ok up to 30-35MW.

MV aks to have access to DFS/Cernbox as a mean to start sharing the data. They are saving data locally at the moment.
Right now, the crate is standalone and not connected. MV asks for the possibility of having a virtual machine with Labview (they have student version). NClH asks difference between Labview student and normal license, maybe some packages missing, not sure. MV are concerned about version of Labview (2018). Agreed that they should maintain it for compatibility, reason for virtual machine.

Regarding the contract, NClH comments that apparently she can do it on her own. She will ask again legal service during this week.

Xbox 2- Update: Now reaching 46MW (around 80MV/m)

PAA corrected plots of BD localization to show full filling time of structure.

PAA confirmed that weird BD apparently coming from wrong acquisition were real BD, due to flag in PKR. NClH comments that usually when going up in power they had to adjust threshold for PKR according to % of peak power.

WW and AG concerned about BDR being very high. PAA confirms limit is 5x10-5, but it could be decreased.  WW says maybe it’s time to run flat for a while. PAA confirms she did it, and it decreased BDR, will do it again during this week.

Comment on some weird BD pulses. Nothing through structure, high reflected power towards klystron. WW and AG comment maybe in PC going on, some multipackting. They also observe previous pulses not detected as BD pulses have high PKR. Could be dangerous to have klystron pulse but nothing to the structure. PAA to calculate integral of power.

NClH says to check pulses and lower threshold for PKR to be sure. IS asks about vacuum level at this time, maybe plot them together to better correlate, could be some gas in cavities (not completely recovered vacuum). Also, IS notices that from pulse to pulse, pulse moves, so cavities are detuned.  

Also, compressed pulse with sometimes factor 2 sometimes factor 4. NClH remembers Marca saying we were losing half the power. AG says we should see some heating.

Action: For next week: running flat at 45MW. See if BDR decreases.

Xbox2- integration of BOC. PAA asks about this, not having enough channels for putting HPDC after BOC. All agree to leave it as it is, no signals between output of klystron and structure.

Xbox3- HEK. Flow meters have arrived now, just waiting for interfaces to connect the tubes.

Surface temperature sensors have been installed in collector in and out water pipes, and correctly ensured isolation and thermal contact. IS says now we may be able to calibrate with this. WW aks about signal structure of temperature sensors, period of 15min. NClH confirms this is due to water station valve response speed. Used to be worse, this was the improvement. Still, WW says this is big source of uncertainty for calorimetry. We need to be careful to align signals before substracting them.

WW suggests to have calibrated heater. For the moment, we will work like this, IS says should be enough for the accuracy we need. NClH says we have power measurements, calorimetry measurements, and high voltage measurements, that we can correlate.

Finally, PAA has finally correctly trigger the signal analyser to see pulse from driver.


Action: PAA to start installation of temperature probes and flow meters in loads. RB will do plot and analysis of temperature measurements of collector for next week.

There are minutes attached to this event. Show them.
    • 10:00 AM 10:30 AM
      Xboxes report 30m
      Speakers: Mr Lee Millar (CERN), Mareike Wendelmuth (Technische Univ. Hamburg Harburg (DE)), Marçà Boronat (CERN), Dr Matteo Volpi, Paz Alonso Arias (CERN), Xiaowei Wu